OpenStudy (anonymous):

4c2 - 12c + 9 factor completely

OpenStudy (anonymous):

4c^2

OpenStudy (anonymous):

(2c-3)^2 http://www.wolframalpha.com/input/?i=factor+4c2+-+12c+%2B+9

OpenStudy (anonymous):

or i meant (2c-3)^2=0

OpenStudy (anonymous):

can u show me how u got tht?

OpenStudy (anonymous):

4c^2 - 12c + 9 = (2c-3)(2c-3) = (2x-3)^2

OpenStudy (anonymous):

i use wolframalpha but i have to show the work :/ so thanx guyz